A company's profit for the second half year is $150,000 greater than the profit for the first half year. The total annual profit is $850,000. What is the profit for the first half yaer?
Found 2 solutions by amoresroy, SUNNIE630:
Answer by amoresroy(361)

Let x = profit for 1st half in dollars
x + x + 150,000 = 850,000
2x = 850,000 - 150,000
2x = 700,000
x = 350,000

Answer by SUNNIE630(1)

P = Profit
P+P + 150,000 = 850,000
2P = 850,000 - 150,000
2P = 700,000

Divide 700,000 by 2 and you will get 350,000 so that means P= 350,000
Answer is 350,000 for the first half year